# Towards Testing and Evaluating Vision-Language-Action Models for Robotic Manipulation: An Empirical Study

### v1

@@ Abstract.
Multi-modal foundation models and generative AI have demonstrated promising capabilities in applications across various domains. Recently, Vision-language-action (VLA) models have attracted much attention regarding their potential to advance robotic manipulation. Despite the end-to-end perception-control loop offered by the VLA models, there is a lack of comprehensive understanding of the capabilities of such models and an automated testing platform to reveal their robustness and reliability across different robotic manipulation scenarios. To address these challenges, in this work, we present VLATest, a testing framework that automatically generates diverse robotic manipulation scenes to assess the performance of VLA models from various perspectives. Large-scale experiments are considered, including eight VLA models, four types of manipulation tasks and over 18,604 testing scenes. The experimental results show that existing VAL models still lack imperative robustness for practical applications. Specifically, the performance of VLA models can be significantly affected by several factors from the operation environments such as camera poses, lighting conditions and unseen objects. Our framework and the insights derived from the study are expected to pave the way for more advanced and reliable VLA-enabled robotic manipulation systems in practice.
@@
$c$s: Software and its engineering Software defect analysisccs: General and reference Empirical studiesccs: Computer systems organization Robotics
@@ 1. Introduction
Meanwhile, the emergence of foundation models, such as large language models (LLMs) and vision-language models (VLMs), has introduced new opportunities for advancing AI-enabled robotic manipulation (Singh et al., 2023; Ding et al., 2023). In particular, LLMs and VLMs have demonstrated promising capabilities to participate in the loop of robotics system development and operation to enhance corresponding reasoning, perception, and task-planning abilities  (Huang et al., 2023a; Liang et al., 2023b; Driess et al., 2023). In contrast to the general purpose-oriented LLMs and VLMs, vision-language-action (VLA) models were developed exclusively to generate robot actions for manipulation based on visual observations from cameras and task instructions provided in users’ natural language input. Techniques such as reinforcement learning from human feedback (RLHF) (Bai et al., 2022) further enhance these models’ ability to interpret human intent from natural language commands. As a result, it is now possible to let robotics perform manipulation tasks simply by prompting natural language commands into the VLA models. Additionally, the large-scale pre-training nature of such foundation models makes it possible to adapt to a diverse range of downstream manipulation tasks without task- and environment-specific niche design. A recent study demonstrates that the state-of-the-art (SOTA) VLA model, RT-2 (Brohan et al., 2023), can be extended to complex tasks such as reasoning, symbol understanding, and human recognition—tasks that were unseen during training (Brohan et al., 2023).
@@
However, the data-driven nature of VLA models makes them difficult to interpret, which hinders the safety and applicability of VLA models to more robotics applications in practice. To enhance the reliability and trustworthiness of VLA models for robotic manipulation, the development of quality assurance techniques such as testing, debugging, and repairing has become an urgent need in both industrial and academic communities. In particular, comprehensive and adequate testing is essential for throughout assessing a VLA model’s capabilities and limitations. Unfortunately, current VLA models are typically evaluated using only a limited set of hand-crafted scenes, where the comprehensiveness and effectiveness of such scenes are often inadequate, which subsequently leaves the general performance and behavior characteristics of VLA models largely unexplored. Moreover, unlike conventional LLMs, which only process Natural Language (NL) text inputs, VLA models are empowered by the multimodality perception ability; therefore, when deploying VLA models in real-world scenarios, visual factors such as obstacles, lighting conditions, and camera angles may significantly impact their performance. Despite this, there is currently a lack of deep understanding regarding the robustness of VLA models against the variations among such environmental factors.
@@
To address this gap, we propose VLATest, one of the first testing frameworks specifically designed for evaluating VLA models in robotic manipulation. We introduce a set of ten testing operators and design a generation-based fuzzing framework that automatically produces testing scenes to assess the performance and robustness of VLA models in specific robotic manipulation tasks. VLATest is implemented within the Maniskill2 simulation environment (Gu et al., 2023). By leveraging VLATest, we facilitate efficient testing and enable comprehensive evaluation of VLA models under varying conditions. Based on VLATest, we performed a large-scale empirical study to evaluate the overall performance and robustness of popular VLA models. A set of research questions are proposed to examine key factors in reliable robotic manipulation with VLA: basic performance, task complexity, perception robustness, and OOD (out-of-distribution) robustness. Specifically, we investigated the following five research questions:
@@
    How does the number of obstacles affect a VLA model’s performance?
@@
To investigate these research questions, we generated 18,604 testing scenes across four different robotic manipulation tasks. We selected seven popular publicly available VLA models: RT-1-1k, RT-58k, RT-400k (Brohan et al., 2022), RT-1-X (Padalkar et al., 2023), Octo-small, Octo-base (Team et al., 2024), and OpenVLA-7b (Kim et al., 2024). The experiments included a total of 78,604 rounds of simulation execution, which took more than 400 GPU hours.
Our analysis shows that current VLA models exhibit subpar performance across the four robotic manipulation tasks studied. As the number of obstacles increases, it becomes more challenging for these models to accurately locate and manipulate the correct objects. Additionally, we observed that the subject VLA models lack robustness when subjected to changes in lighting conditions and camera angles, performing much worse compared to the default settings. Meanwhile, we found that VLA models with large-scale pre-training generally show better robustness against such changes. Finally, we found that these VLA models struggle significantly with unseen objects, exhibiting a significant drop in performance compared to manipulating seen objects. Our findings provide practical guidelines for developers working with VLA models on specific robotic manipulation tasks and underscore the need for more advanced VLA models, as well as novel quality assurance techniques to enhance the robustness and reliability of the VLA models.
@@
  • A testing framework. We designed and implemented a generation-based fuzzing framework, VLATest, to test VLA models by incorporating various operators in robotic manipulation testing scenes.
  • An empirical study. Based on VLATest, we conducted a large-scale empirical study to evaluate the performance and robustness of seven popular VLA models across four robotic manipulation tasks under various conditions, including variations in the number of obstacles, lighting conditions, camera poses, and the manipulation of both seen and unseen objects.
@@ 2. Background: Vision-Language-Action Models
Figure 2. Operators used for testing VLA models in VLATest.
@@ 2.1. VLA Model for Robotic Manipulation
Fig. 1 illustrates the common architecture of a VLA model. Given a natural language input $T$ and an image input $I_{1}$ at the first timestamp, the VLA model’s natural language tokenizer and visual encoder first project $T$ and $I_{1}$ into sets of tokens, $\mathbf{T}=\{t_{1},\dots,t_{m}\}$ and $\mathbf{I}_{1}=\{i_{11},\dots,i_{1n}\}$, respectively. These tokens are then concatenated and fed into a transformer model to predict action token(s) $\mathbf{A}_{1}$. An action head layer then de-tokenizes $\mathbf{A}_{1}$ into robot action values $\Delta_{x1},\Delta_{\theta 1},\Delta_{grip1}]$, where $\Delta_{x1}\in\mathbb{R}^{3}$, $\Delta_{\theta 1}\in\mathbb{R}^{3}$, and $\Delta_{grip1}\in\mathbb{R}^{1}$ denote the translation ($x,y,z$ movement), rotation, and gripper actions of the end-effector that should be performed after the observation, respectively.
@@ 2.2. Training and Evaluation of the VLA Model
Training. There are typically two approaches to training a VLA model: (1) training from scratch and (2) fine-tuning a general-purpose vision-language model (VLM). Training from scratch involves building a VLA model directly on robot demonstration data. This approach is commonly applied for models with relatively small-scaled architectures and limited computational resources, such as RT-1 (Brohan et al., 2022) and Octo (Team et al., 2024), which typically have fewer than 100 million parameters. By contrast, fine-tuning from an existing VLM leverages the flexibility of re-using a larger model (more than 1 billion parameters), such as Llava (Liu et al., 2024a; Kim et al., 2024), which has been pre-trained on massive amounts of image and text data from diverse domains. The large-scale pre-training of these VLMs may enhance the generalizability of the fine-tuned VLA models, particularly when manipulating unseen objects and tasks.
Evaluation. A VLA model is usually evaluated by measuring its performance on specific skills, such as picking up an object in a scene. To evaluate a particular skill, developers must first create an adequate testing scene, which involves configuring target objects, obstacles, and environmental factors like lighting conditions. Additionally, a text prompt should be meticulously crafted before executing the VLA model to perform the manipulation task. To assess the model’s performance, developers usually design a set of evaluation metrics. For instance, when performing the task “picking up an object”, the metrics include: (1) whether the robot grasps the correct object, (2) whether the object is successfully lifted, and (3) whether the robot can sustain lifting the object for a short period. The execution of the VLA model and the robot can occur in either a simulated or real-world environment. In real-world scenarios, manual labeling is required to compute these metrics, whereas, in simulated environments, the evaluation process can be automated.
@@ 3.1. Operators
We consider four categories of testing operators in VLATest as shown in Fig. 2, resulting in a total of ten testing operators.
@@
  • Obstacles. Different from target object(s), obstacles are typically not part of the task, and the robot is expected to avoid contacting these objects during manipulation. In addition to the three operators for target object(s), we also consider the number of obstacles as a distinct operator in the obstacles category.
  • Lighting. The lighting condition affects the rendering of images captured by the camera, thereby influencing the visual input of the VLA models. In VLATest, we consider lighting intensity as one of the operators.
@@ 3.2. Testing Scene Generation
We summarize the algorithm for generating a testing scene in Algorithm 1. Intuitively, there are four distinct steps: (1) sampling semantically valid target object(s), (2) sampling obstacles (optional), (3) mutating lighting conditions (optional), and (4) mutating camera poses (optional).
Sampling semantically valid target object(s). For a robotic manipulation task, we first randomly sample target object(s) without replacement from the object database $\mathcal{O}$ (Lines 5-9). We also check if the selected objects are semantically valid according to the task requirements (Line 4). For example, in a task requiring object A to be placed over object B, if object B is a ball, object A can never be stably placed on it. Such cases are considered semantically invalid. After selecting the target object(s), we proceed to sample their positions and orientations (Line 7). In our empirical study, we used a random sampler for pose selection. If there are multiple target objects, we ensure they remain far enough apart (>safe_dist>safe\_dist) to avoid generating invalid testing scenes. For instance, in the previous example, if object A is already placed on object B during the sampling of their positions, the testing scene becomes invalid.
Input: an object database $\mathcal{O}$, the robotic manipulation task $T$, the number of target objects $N_{target}$, the number of obstacles $N_{obstacles}$, the lighting mutation option lighting_flaglighting\_flag, and the camera mutation option camera_flagcamera\_flag, the minimum distance between objects safe_distsafe\_dist
@@ 1 $\mathcal{O}~\leftarrow~\mathsf{random\_shuffle}(\mathcal{O}$);
2 $O_{target},O_{obstacles}~\leftarrow~\emptyset,\emptyset$;
[…]
@@ 8     $O_{target}~\leftarrow~O_{target}\bigcup(obj,pos,ori)$;
16     $pos,ori~\leftarrow~\mathsf{pose\_sampler}(O_{target}\bigcup O_{obstacles},$ safe_dist)safe\_dist);
17     $O_{obstacles}~\leftarrow~O_{target}\bigcup(obj,pos,ori)$;
@@
22if lighting_flaglighting\_flag then
@@ 23  $intensity~\leftarrow~\mathsf{mutate\_lighting}(intensity)$;
26if camera_flagcamera\_flag then
@@ 27  $cam_{pos},cam_{ori}~\leftarrow~\mathsf{mutate\_camera}(cam_{pos},$ $cam_{ori})$;
30return $\{O_{target},O_{obstacles},intensity,(cam_{pos},cam_{ori})\}$;
@@
Sampling obstacles. If the number of obstacles $N_{obstacles}$ is greater than 0 (Line 11), we further randomly sample obstacle objects from $\mathcal{O}$ without replacement (Lines 12-16). For each obstacle, we follow the same procedure as used for sampling the target objects to assign their position and orientation (Line 14). In our empirical study, we utilized a random sampler to generate the position and orientation for each obstacle.

### v2

# VLATest: Testing and Evaluating Vision-Language-Action Models for Robotic Manipulation

@@ Abstract.
The rapid advancement of generative AI and multi-modal foundation models has shown significant potential in advancing robotic manipulation. Vision-language-action (VLA) models, in particular, have emerged as a promising approach for visuomotor control by leveraging large-scale vision-language data and robot demonstrations. However, current VLA models are typically evaluated using a limited set of hand-crafted scenes, leaving their general performance and robustness in diverse scenarios largely unexplored. To address this gap, we present VLATest, a fuzzing framework designed to generate robotic manipulation scenes for testing VLA models. Based on VLATest, we conducted an empirical study to assess the performance of seven representative VLA models. Our study results revealed that current VLA models lack the robustness necessary for practical deployment. Additionally, we investigated the impact of various factors, including the number of confounding objects, lighting conditions, camera poses, unseen objects, and task instruction mutations, on the VLA model’s performance. Our findings highlight the limitations of existing VLA models, emphasizing the need for further research to develop reliable and trustworthy VLA applications.
@@
copyright: ccdoi: 10.1145/3729343journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE073publicationmonth: 7ccs: Software and its engineering Software defect analysisccs: General and reference Empirical studiesccs: Computer systems organization Robotics
@@ 1. Introduction
Meanwhile, the emergence of foundation models, such as large language models (LLMs) and vision-language models (VLMs), has introduced new opportunities for advancing AI-enabled robotic manipulation (Singh et al., 2023; Ding et al., 2023). In particular, LLMs and VLMs have demonstrated promising capabilities to participate in the loop of robotics system development and operation to enhance corresponding reasoning, perception, and task-planning abilities (Huang et al., 2023; Liang et al., 2023b; Driess et al., 2023). In contrast to the general purpose-oriented LLMs and VLMs, vision-language-action (VLA) models were developed exclusively to generate robot actions for manipulation based on visual observations from cameras and task instructions provided by users’ natural language input. Techniques such as reinforcement learning from human feedback (RLHF) (Bai et al., 2022) further enhance these models’ ability to interpret human intent from natural language commands. As a result, it is now possible to let robotics perform manipulation tasks simply by prompting the VLA models. Additionally, the large-scale pre-training nature of such foundation models makes it possible to adapt to a diverse range of downstream manipulation tasks without task- and environment-specific niche design. A recent study demonstrates that the state-of-the-art (SOTA) VLA model, RT-2 (Brohan et al., 2023), can be extended to complex tasks such as symbol understanding and human recognition that were unseen during training (Brohan et al., 2023).
@@
However, the data-driven nature of VLA models makes them difficult to interpret, which hinders the safety and applicability of VLA models to more robotics applications in practice. To enhance the reliability and trustworthiness of VLA models for robotic manipulation, the development of quality assurance techniques such as testing, debugging, and repairing has become an urgent need in both industrial and academic communities. In particular, comprehensive and adequate testing is essential for assessing a VLA model’s capabilities and limitations. Unfortunately, current VLA models are typically evaluated using only a limited set of hand-crafted scenes, where the comprehensiveness and effectiveness of such scenes are often inadequate, which subsequently leaves the general performance and behavior characteristics of VLA models largely unexplored. Moreover, unlike conventional LLMs, which only process natural language text inputs, VLA models are empowered by the multimodality perception ability; therefore, when deploying VLA models in real-world scenarios, visual factors such as confounding objects, lighting conditions, and camera angles may significantly impact their performance. Despite this, there is currently a lack of deep understanding regarding the robustness of VLA models against such environmental factors.
@@
To address this gap, we conducted a large-scale empirical study to evaluate the overall performance and robustness of popular VLA models. We first propose VLATest, one of the earliest testing frameworks specifically designed for evaluating VLA models in robotic manipulation. We introduce a set of ten testing operators and design a generation-based fuzzing framework that automatically produces testing scenes to assess the performance and robustness of VLA models in specific robotic manipulation tasks. VLATest is implemented within the Maniskill2 simulation environment (Gu et al., 2023). By leveraging VLATest, we aim to efficiently identify potential bugs in using VLA models for robotic manipulation, enabling comprehensive evaluation of VLA models under varying conditions. We designed a set of research questions to examine key factors in reliable robotic manipulation with VLA: basic performance, task complexity, perception robustness, OOD (out-of-distribution) robustness, and language model robustness. Specifically, we investigated the following six research questions:
@@
    How does the number of confounding objects affect a VLA model’s performance?
@@
  • RQ6
    How robust do VLA models perform against task instruction mutations?
To investigate these research questions, we generated 18,604 testing scenes across four different robotic manipulation tasks. We selected seven popular publicly available VLA models: RT-1-1k, RT-58k, RT-400k (Brohan et al., 2022), RT-1-X (Padalkar et al., 2023), Octo-small, Octo-base (Team et al., 2024), and OpenVLA-7b (Kim et al., 2024). The experiments included a total of 78,604 rounds of simulation execution, which took more than 580 GPU hours.
Our analysis shows that current VLA models exhibit subpar performance across the four robotic manipulation tasks studied. As the number of confounding objects increases, it becomes more challenging for these models to accurately locate and manipulate the correct objects. Additionally, we observed that the subject VLA models lack robustness when subjected to changes in lighting conditions and camera angles, performing much worse compared to the default settings. Meanwhile, we found that VLA models with large-scale pre-training generally show better robustness against such changes. Finally, we found that these VLA models struggle significantly with unseen objects, exhibiting a significant drop in performance compared to manipulating seen objects. Current VLA models also often fell short in understanding the same task instruction paraphrased in different ways. Our findings provide practical guidelines for developers working with VLA models on specific robotic manipulation tasks and underscore the need for more advanced VLA models, as well as novel quality assurance techniques to enhance the robustness and reliability of the VLA models.
@@
  • An empirical study. We conducted a large-scale empirical study to evaluate the performance and robustness of seven popular VLA models across four robotic manipulation tasks under various conditions, including variations in the number of confounding objects, lighting conditions, camera poses, and the manipulation of both seen and unseen objects.
  • A testing framework. To conduct our empirical study, we designed and implemented a generation-based fuzzing framework, VLATest, to test VLA models by incorporating various operators in robotic manipulation tasks.
@@
  • Artifacts. Our artifacts, including the replication packages and the generated testing scenes, are available on a GitHub repository: https://github.com/ma-labo/VLATest.
@@ 2.1. VLA Model for Robotic Manipulation
Fig. 1 illustrates the common architecture of a VLA model. Given a natural language input $T$ and an image input $I_{1}$ at the first timestamp, the VLA model’s natural language tokenizer and visual encoder first project $T$ and $I_{1}$ into sets of tokens, $\mathbf{T}=\{t_{1},\dots,t_{m}\}$ and $\mathbf{I}_{1}=\{i_{11},\dots,i_{1n}\}$, respectively. These tokens are then concatenated and fed into a transformer model to predict action token(s) $\mathbf{A}_{1}$. An action head layer then de-tokenizes $\mathbf{A}_{1}$ into robot action values $[\Delta_{x1},\Delta_{\theta 1},\Delta_{grip1}]$, where $\Delta_{x1}\in\mathbb{R}^{3}$, $\Delta_{\theta 1}\in\mathbb{R}^{3}$, and $\Delta_{grip1}\in\mathbb{R}^{1}$ denote the translation ($x,y,z$ movement), rotation, and gripper actions of the end-effector that should be performed after the observation, respectively.
@@ 2.2. Training and Evaluation of the VLA Model
Training. There are typically two approaches to training a VLA model: (1) training from scratch and (2) fine-tuning a general-purpose VLM. Training from scratch involves building a VLA model directly on robot demonstration data. This approach is commonly applied for models with relatively small-scale architectures and limited computational resources, such as RT-1 (Brohan et al., 2022) and Octo (Team et al., 2024), which typically have fewer than 100 million parameters. By contrast, fine-tuning from an existing VLM leverages the flexibility of re-using a larger model (more than 1 billion parameters), such as Llava (Liu et al., 2024a; Kim et al., 2024), which has been pre-trained on massive amounts of image and text data from diverse domains. The large-scale pre-training may enhance the generalizability of the fine-tuned VLA models, particularly when manipulating unseen objects and tasks.
Evaluation. A VLA model is usually evaluated by measuring its performance on specific skills, such as picking up an object in a scene. To evaluate a particular skill, developers must first create an adequate testing scene, which involves configuring target objects, confounding objects, and environmental factors such as lighting conditions. Additionally, a text prompt should be meticulously crafted before performing the manipulation task. To assess the model’s performance, developers usually design a set of evaluation metrics. For instance, when performing the task “picking up an object”, the metrics include: (1) whether the robot grasps the correct object, (2) whether the object is successfully lifted, and (3) whether the robot can sustain lifting the object for a short period. The execution of the VLA model and the robot can occur in either a simulated or real-world environment. In real-world scenarios, manual labeling is required to compute these metrics, whereas in simulated environments, the evaluation process can be automated.
@@ 3.1. Operators
We consider four categories of testing operators in VLATest as shown in Fig. 2, resulting in a total of ten testing operators. We choose these operators because they are the most essential ones in the configuration of a robotic manipulation task (Siciliano, 2008; Murray et al., 2017). Thus, VLATest can be easily adapted for a wide range of tasks. Additionally, by only changing these operators, our generated scenes will not change the intrinsic nature of a manipulation task. This ensures that all test cases will not require the VLA models to perform actions that were never performed during training.
Figure 2. Operators used for testing VLA models in VLATest.
@@
  • Confounding objects. Different from target object(s), confounding objects are typically not part of the task, and the robot is expected to avoid contacting these objects during manipulation. In addition to the three operators for target object(s), we also consider the number of confounding objects as a distinct operator in the confounding objects category.
  • Lighting. The lighting condition affects the rendering of images captured by the camera, thereby influencing the visual input of the VLA models. We consider lighting intensity as one operator.
@@ 3.2. Testing Scene Generation
We summarize the algorithm for generating a testing scene in Algorithm 1. Intuitively, there are four distinct steps: (1) sampling semantically valid target object(s), (2) sampling confounding objects (optional), (3) mutating lighting conditions (optional), and (4) mutating camera poses (optional).
Sampling semantically valid target object(s). For a robotic manipulation task, we first randomly sample target object(s) without replacement from the object database $\mathcal{O}$ (Lines 5-9). We also check if the selected objects are semantically valid according to the task requirements (Line 4). For instance, in a task requiring object A to be placed over object B (Task 3), it is possible to put an apple on a plate, but it is impractical to put an apple on a ball, since placing an apple on a ball stably could be very challenging. Similarly, for Task 4, it is impossible to put an apple inside a Coke can. We consider scenes that can not be successfully manipulated by the robots to be semantically invalid. We manually created a list of invalid objects for tasks 3 and 4 to check if a generated test scene is semantically valid. After selecting the target object(s), we proceed to sample their positions and orientations (Line 7). In our empirical study, we used a random sampler for pose selection. If there are multiple target objects, we ensure they remain far enough apart ($>safe\_dist$) to avoid generating invalid testing scenes. For instance, in the previous example, if object A is already placed on object B during the sampling of their positions, the testing scene becomes invalid.
Input: an object database $\mathcal{O}$, the robotic manipulation task $T$, the number of target objects $N_{target}$, the number of confounding objects $N_{confound}$, the lighting mutation option $lighting\_flag$, and the camera mutation option $camera\_flag$, the minimum distance between objects $safe\_dist$
@@ 1 $\mathcal{O}~{}\leftarrow~{}\mathsf{random\_shuffle}(\mathcal{O}$);
2 $O_{target},O_{confound}~{}\leftarrow~{}\emptyset,\emptyset$;
3 $intensity,cam_{pos},cam_{ori}~{}\leftarrow~{}$default values;
@@ 8             $O_{target}~{}\leftarrow~{}O_{target}\bigcup(obj,pos,ori)$;
16             $pos,ori~{}\leftarrow~{}\mathsf{pose\_sampler}(O_{target}\bigcup O_{confound},$ $safe\_dist)$;
17             $O_{confound}~{}\leftarrow~{}O_{target}\bigcup(obj,pos,ori)$;
@@
22if $lighting\_flag$ then
@@ 23       $intensity~{}\leftarrow~{}\mathsf{mutate\_lighting}(intensity)$;
26if $camera\_flag$ then
@@ 27       $cam_{pos},cam_{ori}~{}\leftarrow~{}\mathsf{mutate\_camera}(cam_{pos},$ $cam_{ori})$;
30return $\{O_{target},O_{confound},intensity,(cam_{pos},cam_{ori})\}$;
@@
Sampling confounding objects. If the number of confounding objects $N_{confound}$ is greater than 0 (Line 11), we further randomly sample confounding objects from $\mathcal{O}$ without replacement (Lines 12-16). Note that we consider similar objects with different semantics to be different. For instance, a Coke can is considered different from a Pepsi can. Therefore, having both of them as confounding objects in one scene is possible. For each confounding object, we follow the same procedure as used for sampling the target objects to assign their position and orientation (Line 14). In our empirical study, we utilized a random sampler to generate positions and orientations.
@@
After generating each scene, we compared its configuration with existing ones to prevent duplication. Throughout our experiments, no identical test scenes were generated.